OpenStudy (anonymous):

differentiate. [ln(x)]/[1+ln(2x)] please show steps

OpenStudy (kainui):

Quotient rule says d/dx(f(x)/g(x)) = (f'(x)g(x)-g'(x)f(x))/(g(x))^2 Or more simply, (the derivative of the top times the bottom) minus (the derivative of the bottom times the top) all over the bottom squared. (1/x) = derivative of ln(x) (1/x)(1+ln2x)-(1/2x)(lnx) --------------------- (1+ln2x)^2

OpenStudy (anonymous):

i dont use chain rule on the bottom?

OpenStudy (kainui):

The bottom is just squared, you don't take the derivative of the bottom.

OpenStudy (anonymous):

oh, right. isnt the derivative of 1+ln2x = 2/2x?

OpenStudy (kainui):

Yeah, you're right the derivative of ln(2x) is 2/2x due to the chain rule, whoops.

OpenStudy (anonymous):

alrite cool. thamks dude.